Roger can run one mile in 16 minutes. Jeff can run one mile in 14 minutes. If Jeff gives Roger a 1 minute head start, how long will it take before Jeff catches up to roger? How far will each have run?
Answer by stanbon(57369)

Roger can run one mile in 16 minutes. Jeff can run one mile in 14 minutes. If Jeff gives Roger a 1 minute head start, how long will it take before Jeff catches up to roger? How far will each have run?
--------------
Roger DATA:
rate = 1/16 mile/minute ; time = x minutes distance = rt = x/16 miles
------
Jeff DATA:
rate = 1/14 mile/minute ; time = x-1 minutes ; distance = rt = (x-1)/14 miles
=============================================
Equation:
distance = distance
x/16 = (x-1)/14
14x = 16x-16
2x = 16
x = 8 minutes (Roger time)
Roger will run rt = (1/16)8 = 1/2 mile
x-1 = 7 minutes (Jeff time)
Jeff will run rt = (1/14)7 = 1/2 mile
================================
Cheers,
Stan H.
==============